OpenStudy (anonymous):

*~*Please Help*~* Solve –3(2x – 4) = 24

OpenStudy (anonymous):

x=6 i believe.

OpenStudy (ajprincess):

–3(2x – 4) = 24 -6x+12=24 1.Add 6x both sides. 2.Subtract 24 both sides. 3.Divide by 6 both sides.

OpenStudy (ajprincess):

Can u do?

OpenStudy (anonymous):

Yeah I can thank you! c:

OpenStudy (ajprincess):

yw.:)